OpenStudy (anonymous):

What are the exact solutions of x2 - 5x - 1 = 0?

OpenStudy (anonymous):

first you attempt to factor if that doesnt work use the quadratic equation

OpenStudy (anonymous):

platinum is the x squared?

OpenStudy (anonymous):

Use quadratic formula: a = 1, b = -5, c = -1 x = [-b±√(b²-4ac)]/(2a) x = {-(-5)±√[(-5)²-4(1)(-1)]}/[2(1)] x = {5±√[(25)+(4)]}/2 x = [5±√(29)]/2 Solutions: x = [5±√(29)]/2